A lead-acid battery uses a redox reaction in whichlead(0) and lead(IV) are both converted to lead(II). This reaction is facilitated by the presence of sulfuric acid, H2SO4, as shown by the reaction

Pb+PbO2+2H2SO4→2PbSO4+2H2O

Suppose that a fully charged lead-acid battery contains 1.50 L of 5.00 M H2SO4. What will be the concentration of H2SO4 in the battery after 4.00 A of current is drawn from the battery for 5.50 hours ?
Answer #1

coulombs= 5.5*60*60* 4 =79200 columbs

This has to be divided by 96500 to get numbef of electrons

number of eletrons =79200/96500=0.821 moles

sulfuric acid consumed =0.821*2= 1.642 moles ( as per the reaction)

moles of sulfuric acid initially present = 5*1.5 = 7.5 moles

moels remaining = 7.5- 1.642 =5.858 moles

Molar concentration = 5.858/1.5 =3.9M